# CLARIFID: Improving Radiology Report Generation by Reinforcing Clinically Accurate Impressions and Enforcing Detailed Findings

### v1

@@ Introduction
Table 1 contrasts our model with recent state-of-the-art systems; R2Gen(Chen et al. 2022b), R2GenCMN(Chen et al. 2022a), KiUT(Huang, Zhang, and Zhang 2023), RGRG(Tanida et al. 2023), METransformer(Wang et al. 2023b), R2GenGPT(Wang et al. 2023c), ORGAN(Hou et al. 2023), COMG(Gu et al. 2024), EKAGen(Bu et al. 2024), MLRG(Liu et al. 2025), PromptMRG(Jin et al. 2024). On the MIMIC-CXR full test set we outperform all baselines, including those optimized solely for the Findings section. Our method attains the best Impression CE (P 0.458 / R 0.540 / F1 0.496) and the highest Findings CE recall (0.506) and F1 (0.530), evidencing superior recovery of clinically pertinent details. Trained only on MIMIC-CXR, the same model also leads on IU X-ray (Impression CE F1 0.638) without any domain adaptation, confirming strong cross-dataset generalization. Together, these gains underline the clinical reliability and robustness of our joint Findings-and-Impression optimization strategy.
Ablation Study
Table 2 evaluates the impact of each component on MIMIC-CXR Impression. Adding reinforcement learning (+RL) that maximizes ”Impression” F1 uplifts both ”Findings” and ”Impression” scores, showing that summary-level rewards inject clinically useful signals into the entire reasoning pipeline. Introducing the forced generation (+FG) raises F1 for both sections, supporting our key claim that radiology report generation is inherently sequential—accurate ”Findings” naturally set the stage for concise ”Impression”. Finally, the Best-of-N (+BoN) provides the strongest overall performance, raising “Findings” F1 to 0.530, “Impression” F1 to 0.496, and Meteor to 0.179.
Section-Wise Effects of Reinforcement Learning
@@
To obtain a more detailed analysis of the effect of reinforcement learning (RL), we tracked section-level F1 score during reinforcement-learning training. Figure 4 presents these scores on the MIMIC-CXR Impression test set across 1,275 RL iterations. The Impression F1 exhibits a consistent upward trajectory from $0.416$ to $0.479$, while the Findings F1 likewise advances from $0.493$ to $0.507$ over the same interval. These parallel improvements indicate that an Impression-based reward not only strengthens the Impression section itself but also yields measurable gains in the Findings section.
Qualitative result
To further demonstrate the effectiveness of our approach, we conducted a qualitative evaluation on the MIMIC-CXR test set. For each disease finding, we overlaid color highlights on text segments where the generated report semantically matched the ground-truth reference. As shown in Figure 3, increasing the decoding budget at test time leads to a higher degree of overlap between generated and reference findings, indicating that our controlled budget-forcing strategy yields more clinically faithful reports. Furthermore, key findings identified in the Findings section are reiterated in the Impression section, ensuring that core diagnostic information is consistently emphasized.
Conclusion
We have presented CLARIFID, a unified radiology-report generation framework that explicitly mirrors the radiologist’s two-step reasoning process—first documenting ”Findings”, then synthesizing an ”Impression”. By directly optimizing the clinical correctness of the Impression section and enforcing reasoning-aware decoding, our approach aligns both training and inference with real-world diagnostic practice. This synergy delivers state-of-the-art performance across benchmarks, boosting Impression-level clinical efficacy on MIMIC-CXR and, without any domain tailoring, outperforming Impression-level F1 on IU X-ray which is the strong evidence of robust cross-dataset generalization.
Limitations
Although CLARIFID markedly improves clinical fidelity, several limitations remain. First, prioritizing impression-level accuracy can narrow linguistic diversity, yielding lower BLEU and ROUGE scores than text-centric baselines. Second, our evaluation relies entirely on automatic metrics; blinded reviews by practizing radiologists are still needed to verify real-world utility and expose subtle errors. Future work will seek a better balance between clinical faithfulness and language naturalness, and incorporate expert human assessment to address these issues.
Appendix A Appendix
Network Architecture
Policy Network Architecture
The policy network employs a GPT-2 decoder with 30 transformer layers, hidden size 768, and 12 attention heads. It processes the padded image representations $\widehat{\mathbf{S}}\in\mathbb{R}^{B\times M_{max}\times D}$ and attention mask $\mathbf{A}\in\{0,1\}^{B\times M_{max}}$ through cross-attention at each transformer layer:
$h_{l}^{text}$ $=\text{SelfAttn}(h_{l-1}^{text})$ (8)
$h_{l}^{cross}$ $=\text{CrossAttn}(h_{l}^{text},\widehat{\mathbf{S}},\text{mask}=\mathbf{A})$ (9)
$h_{l}^{final}$ $=\text{FFN}(h_{l}^{cross})$ (10)
The attention mask $\mathbf{A}$ ensures that for each study $i$, only the first $K_{i}\cdot N$ entries in row $i$ (representing valid visual tokens) influence the generation. This architecture enables clinically accurate reports conditioned on all available views, with the policy network outputting a distribution $P_{\theta}(y_{t}|y_{<t},\widehat{\mathbf{S}},\mathbf{A})$ at each step.
Value Network Architecture
The value network shares the same cross-attention mechanism but uses a lighter configuration with 6 transformer layers, while maintaining the same hidden size and attention heads. It receives the same inputs $(\widehat{\mathbf{S}},\mathbf{A})$ as the policy network and maps the current state to a scalar value:
$V_{\phi}(s_{t})=f_{\phi}(\text{GPT2}_{\phi}(y_{<t},\widehat{\mathbf{S}},\mathbf{A}))$ (11)
where $f_{\phi}$ is a linear projection from the final hidden state to a scalar. This value estimate is crucial during PPO training for computing advantages and during inference for best-of-N impression selection.
Additional Implementation Details
Section-Aware Pretraining
During the section-aware pretraining phase we use a learning rate of $1\times 10^{-4}$ and optimize the image encoder, text decoder, and cross-attention block jointly with the same rate. To improve robustness, input images are randomly rotated within ±$\pm 5^{\circ}$, and at every epoch the order of the findings sentences is shuffled. The policy model (our default generator) employs a GPT-2 backbone with 30 transformer layers, whereas the value model uses a 6-layer GPT-2 backbone. The number of input images is capped at three.
PPO training
In the reinforcement-learning phase we begin with a learning rate of $1\times 10^{-6}$, linearly warm it up to $1\times 10^{-5}$ over the first 300 iterations, and keep it constant until iteration 1275. We set the PPO batch size to 2 and use gradient accumulation step size of  128. During roll-outs the sampling temperature is fixed at 1.0. We adopt reward normalization and 4 rollouts per sample following (Havrilla et al. 2024).
Additional Experiments
@@ Appendix A Appendix
Figure 5: Impact of next-token forcing count on clinical accuracy F1 scores for Findings, Impression, and Full Report evaluated on the MIMIC-CXR Impression test set. Scores are plotted across forcing lengths of 2 to 15 tokens without BoN.
@@
Table 3: Performance at different temperature scheduling settings.
CheXbert Label M-CXR Imp. IU X-ray
% $\mathbf{P}\!\uparrow$ $\mathbf{R}\!\uparrow$ $\mathbf{F1}\!\uparrow$ $\mathbf{P}\!\uparrow$ $\mathbf{R}\!\uparrow$ $\mathbf{F1}\!\uparrow$
Enlarged Cardiomediastinum 9.2 0.130 0.082 0.100 2.4 0.088 0.100 0.093
Cardiomegaly 39.1 0.611 0.680 0.643 14.8 0.594 0.298 0.397
Lung Opacity 37.9 0.531 0.418 0.468 18.5 0.478 0.224 0.305
Lung Lesion 6.7 0.231 0.084 0.123 5.2 0.317 0.119 0.173
Edema 18.0 0.463 0.632 0.535 0.9 0.250 0.167 0.200
Consolidation 5.2 0.190 0.095 0.127 0.9 0.059 0.056 0.057
Pneumonia 4.6 0.188 0.297 0.230 0.7 0.038 0.133 0.060
Atelectasis 24.4 0.395 0.474 0.431 5.6 0.200 0.188 0.194
Pneumothorax 1.9 0.542 0.433 0.481 0.5 0.600 0.273 0.375
Pleural Effusion 29.6 0.655 0.745 0.697 3.2 0.385 0.776 0.515
Pleural Other 3.8 0.219 0.115 0.151 1.8 0.143 0.027 0.045
Fracture 5.8 0.125 0.109 0.116 4.8 0.133 0.100 0.114
Support Devices 33.2 0.688 0.773 0.728 5.3 0.266 0.477 0.342
No Finding 5.1 0.185 0.122 0.147 14.4 0.220 0.320 0.261
Table 4: CE metric of each class on MIMIC-CXR Imp. and IU X-ray
Next-Token Forcing Hyperparameter Study
Temperature Scheduling Hyperparameter Study
To identify an effective division of “creativity” Findings and “conciseness” Impression at inference time, we conduct a small grid‑search over the sampling temperatures $T_{\text{find}}$ and $T_{\text{imp}}$. We explored the four settings as follows.
@@
[…]
Overall, the schedule $\mathbf{(1.0,0.8)}$ provides the strongest balance between detailed Findings and well‑formed Impression sections, and we therefore adopt it as the default in all subsequent experiments.
Class-Wise Clinical Efficacy
Table 4 reports the class-wise ”findings” clinical efficacy metrics of our approach on the MIMIC-CXR Impression and IU X-ray test sets. And we also evaluated macro-averaged, sample-averaged precision(P), recall(R), F1.
IU X-ray Qualitative Result
To further assess the generalizability and clinical reliability of our model, we conducted a qualitative evaluation on the IU X-ray dataset. Figure 6 presents representative examples from both normal and abnormal cases. For each study, we compare the generated report with the ground-truth reference, highlighting disease findings and impression statements.
As shown, our model is able to accurately recover key abnormal findings (e.g., pleural effusion, cardiomegaly) and consistently produces clinically faithful impression summaries. In normal cases, the generated reports avoid hallucinating findings and appropriately reflect the absence of significant abnormalities. These results demonstrate that even without domain-specific fine-tuning on IU X-ray, our model maintains strong clinical consistency and interpretability in both typical and challenging scenarios
Additional MIMIC-CXR Qualitative Result
Figures 7 and 8 provide supplementary qualitative examples from the MIMIC-CXR test set. These cases further demonstrate our model’s ability to generate semantically faithful and clinically accurate reports, even when using different phrasing than the reference. Key findings such as thoracic stability and major disease labels are consistently preserved.
Figure 6: Qualitative evaluation on IU X-ray test set.
@@
Figure 8: Additional MIMIC-CXR dataset test case

### v4

@@ 5.1 Comparison with SOTA
Table 1 contrasts our model with recent state-of-the-art systems; R2Gen Chen et al. (2020), R2GenCMN Chen et al. (2022), KiUT Huang et al. (2023), RGRG Tanida et al. (2023), METransformer Wang et al. (2023b), R2GenGPT Wang et al. (2023c),ORGAN Hou et al. (2023), COMG Gu et al. (2023), EKAGen Bu et al. (2024), MLRG Liu et al. (2025), PromptMRG Jin et al. (2024). On the MIMIC-CXR Impression test set we outperform all baselines, including those optimized solely for the Findings section. Our method attains the best Impression CE (P 0.458 / R 0.540 / F1 0.496) and the highest Findings CE recall (0.557) and F1 (0.530), evidencing superior recovery of clinically pertinent details. Trained only on MIMIC-CXR, our model also leads on IU X-ray (Impression-CE F1 0.638) without domain adaptation. In this split, 39.1% of Findings contain no positive CheXbert labels and 60.6% of Impressions are “No Findings.” Because of this asymmetry, Findings—where labels are scarce—tend to yield more restrained scores, while Impression is heavily influenced by the correct identification of “No Finding,” which dominates its label distribution. Together, these gains underline the clinical reliability and robustness of our joint Findings-and-Impression optimization strategy. Note that MIMIC-CXR serves as a Findings-only benchmark, so a broader set of methods is included. In contrast, MIMIC-CXR Impression and IU X-ray require an explicit Impression for section-aware evaluation, and only a subset of baselines provide such outputs.
5.2 Ablation study
[…]
Figure 4: Section-wise and overall F1 scores on the MIMIC-CXR Impression through RL iteration.
5.3 Section-wise effects of reinforcement learning
@@
To obtain a more detailed analysis of the effect of reinforcement learning (RL), we tracked section-level F1 score during reinforcement learning training. Figure 4 presents these scores on the MIMIC-CXR Impression test set across RL iterations. The Impression F1 increases steadily before saturating at approximately $0.465$, whereas the Findings F1 rises and plateaus near $0.5$ over the same interval. These parallel improvements indicate that an Impression-based reward not only strengthens the Impression section itself but also yields measurable gains in the Findings section.
5.4 Intra-report consistency of Findings and Impression
Table 3: CheXbert intra-report consistency on MIMIC-CXR Impression.
Method $R_{e}$ $P_{e}$
Ground Truth 0.569 0.413
R2GenGPT(Deep) 0.189 0.160
Ours(SL) 0.508 0.367
Ours(SL + RL) 0.616 0.469
Table 4: CheXbert intra-report consistency on IU X-ray.
Method $R_{e}$ $P_{e}$
Ground Truth 0.224 0.211
R2GenGPT(Deep) 0.038 0.027
Ours 0.159 0.132
To verify that the diagnostic Impression is supported by the evidence described in the Findings, we introduce a CheXbert-based intra-report consistency probe that measures alignment between disease-level positives stated in the Impression and those documented in the Findings. Prior work has examined this axis—treating discrepancies between the Findings and the Impression sections as detectable errors Min et al. (2022) and incorporating “impression consistency” into a radiologist-derived evaluation framework Liu et al. (2024b). However, to preserve direct comparability with standard Clinical Efficacy (CE) reporting Chen et al. (2020), we introduce two summary metrics—evidence recall and evidence precision. This direct comparability to CE allows improvements to be interpreted on the same label set and scale widely used across datasets and baselines.
For study condition $c\in\{1,\dots,13\}$ (13 disease labels) and section $s\in\{F,I\}$ (Findings, Impression). We get binarized CheXbert label as in Eq. (5):
$g_{s,c}=\mathbbm{1}[\tilde{g}_{s,c}=\text{positive}]\in\{0,1\}$
Let $L^{+}_{F}=\{c:g_{F,c}=1\}$ and $L^{+}_{I}=\{c:g_{I,c}=1\}$ denote the sets of positive labels from Findings and Impression sections of report, respectively. Evidence recall ($R_{e}$) measures the fraction of positive labels asserted in the Impression that are present as evidence in the Findings, whereas evidence precision ($P_{e}$) measure the fraction of positive labels enumerated in Findings that are carried forward into the Impression.
$R_{e}=\frac{|{L}_{I}^{+}\cap{L}_{F}^{+}|}{\max(1,|{L}_{I}^{+}|)},\quad P_{e}=\frac{|{L}_{I}^{+}\cap{L}_{F}^{+}|}{\max(1,|{L}_{F}^{+}|)}.$
Since the goal of $(R_{e},P_{e})$ is to test whether positive diagnostic claims in the Impression are evidenced in the Findings, we compute them over the 13 disease labels, excluding CheXbert’s “No Findings” label.
Table 3 shows a clear improvement obtained by incorporating reinforcement learning (RL) on MIMIC-CXR Impression ( $R_{e}:0.508\rightarrow 0.616,\;P_{e}:0.367\rightarrow 0.469$ ). These results reinforce that optimizing Impression-level clinical correctness not only improves section-wise CE but also tightens the coupling between sections: a larger fraction of Impression positives are explicitly present in the Findings, and the Findings content reflected in the Impression becomes more focused. Moreover, compared to R2GenGPT Wang et al. (2023c), our supervised baseline (SL) already achieves markedly higher consistency ($\Delta R_{e}{=}{+}0.319$, $\Delta P_{e}{=}{+}0.207$). This pattern suggests that our section-aware training also enhances the linkage between the two sections. On IU X-ray (Table 4), our model—trained only on MIMIC-CXR outperforms R2GenGPT by a wide margin ($R_{e}$: 0.159 vs. 0.038, $P_{e}$: 0.132 vs. 0.027), supporting our claim of better cross-dataset consistency. The remaining gap to ground-truth levels reflects domain/style shift.
5.5 Effects of next-token forcing
@@
Figure 5: Impact of next-token forcing count on clinical accuracy F1 scores for Findings, Impression, and Full Report. Scores are plotted across forcing lengths of 2 to 15 tokens without BoN.
Table 5: Performance at different temperature scheduling settings.
@@
5.6 Effects of temperature scheduling
To identify an effective division of “creativity” Findings and “conciseness” Impression at inference time, we conduct a small grid‑search over the sampling temperatures $T_{\text{find}}(T_{\text{f}})$ and $T_{\text{imp}}(T_{\text{i}})$. We explored the four settings as follows.
@@ 5.6 Effects of temperature scheduling
Table 5 summarizes the outcome. The balanced setting $(1.0,1.0)$ yields the best Findings $\text{F1}=0.536$, whereas the hotter Findings decoder $(1.2,0.8)$ sacrifices precision for recall, resulting in a slightly lower F1. Cooling the Impression decoder to $T_{\text{imp}}=0.8$ while keeping $T_{\text{find}}=1.0$ gives the highest Impression $\text{F1}=0.496$, confirming that lower temperature helps the model compose a clearer clinical summary.
@@
Figure 8: Qualitative evaluation on IU X-ray, The term ‘xxxx’ in the reference report denotes keywords that were removed during the de-identification process
5.7 Classification head performance
To quantify the disease-prediction component used during section-aware pretraining, we evaluate the 14-label classification head on a held-out split. As summarized in Table 6, the model attains micro-averaged Precision = 0.438,Recall = 0.450, and F1 = 0.419. These results indicate that the visual encoder captures clinically informative cues. note that this classification head is used for pretraining/analysis and does not participate in report generation at inference time.
Table 6: Multi-label chest X-ray classification performance of the visual encoder.
Average Precision Recall F1
Micro 0.438 0.450 0.419
5.8 Qualitative results
To demonstrate the clinical faithfulness and generalizability of our model, we conducted a qualitative evaluation on both the MIMIC-CXR and IU X-ray datasets. For our analysis, we overlaid color highlights on text segments where the generated report semantically matched the ground-truth reference, allowing for a direct comparison of clinical content.
5.8.1 MIMIC-CXR results
On the MIMIC-CXR test set, our model consistently generates semantically rich and clinically accurate reports. As shown in Figures 6 and 7, our controlled budget-forcing strategy demonstrates clear benefits; increasing the decoding budget at test time leads to a higher degree of overlap with the reference, yielding more clinically faithful reports. Furthermore, the model successfully preserves key findings (e.g., thoracic stability) and ensures that critical diagnostic information identified in the Findings section is coherently reiterated in the Impression section. This semantic integrity is maintained even when the generated report uses different phrasing than the ground-truth reference.
As shown in figure 9 class-targeted Grad-CAM  Selvaraju et al. (2016) visualizations indicate that the model’s attention is appropriately concentrated on the relevant anatomical regions associated with each diagnostic concept. This suggests that the model does not rely on spurious correlations, but instead focuses on clinically meaningful image regions during the reasoning process that precedes report generation. Overall, the qualitative heatmaps confirm that the visual encoder successfully captures disease-related cues from the appropriate anatomical structures.
Figure 9: Class-targeted Grad-CAM visualizes organ/region-specific heatmaps used by the model when composing the report.
5.8.2 IU X-ray results
To assess the model’s performance in a different domain without specific fine-tuning, we evaluated it on the IU X-ray dataset. The representative examples in Figure 8 showcase the model’s strong generalizability. It accurately identifies key abnormal findings (e.g., pleural effusion, cardiomegaly) in pathological cases while avoiding hallucinated findings in normal cases, appropriately reflecting the absence of significant abnormalities. These results confirm that our model maintains high clinical consistency and interpretability across both typical and challenging scenarios, even on unseen data.
6 Conclusion
We have presented CLARIFID, a unified radiology-report generation framework that explicitly mirrors the radiologist’s two-step reasoning process—first documenting “Findings”, then synthesizing an “Impression”. By directly optimizing the clinical correctness of the Impression section and employing controlled decoding, our approach aligns both training and inference with real-world diagnostic practice. This synergy delivers state-of-the-art performance across benchmarks, boosting Impression-level clinical efficacy on MIMIC-CXR and, without any domain tailoring, achieves the best Impression-level F1 on IU X-ray, providing strong evidence of robust cross-dataset generalization.
Limitations
Although CLARIFID markedly improves clinical fidelity, several limitations remain. First, prioritizing impression-level accuracy can narrow linguistic diversity, yielding lower BLEU and ROUGE scores than text-centric baselines. Second, our evaluation relies entirely on automatic metrics; blinded reviews by practicing radiologists are still needed to verify real-world utility and expose subtle errors. Future work will seek a better balance between clinical faithfulness and language naturalness, and incorporate expert human assessment to address these issues.
CRediT authorship contribution statement
Kyeongkyu Lee: Writing—original draft, Writing—review & editing, Visualization, Validation, Investigation, Conceptualization, Data curation, Software.
Seonghwan Yoon: Writing—original draft, Writing—review & editing, Visualization, Validation, Investigation, Conceptualization, Data curation, Software.
Hongki Lim: Writing—review & editing, Supervision, Conceptualization, Methodology, Project administration, Data curation, Software, Resources.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve the clarity and readability of the manuscript. After using this service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgment
This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-24683103), in part by Korea Basic Science Institute (National Research Facilities and Equipment Center) grant funded by the Ministry of Science and ICT (No. RS-2024-00401899), and in part by Institute of Information & Communications Technology Planning & Evaluation (IITP) under the Leading Generative AI Human Resources Development (IITP-2025-RS-2024-00360227) grant funded by the Korea government (MSIT).
Data Availability
All data used in this work are from publicly available sources.